OpenStudy (hrhsbull):

What is the degree measure of an arc 4 ft. long in a circle of radius 10 ft.?

hero (hero):

x = degree measure y = length of arc \(\dfrac{x}{360} = \dfrac{y}{2 \pi r}\) y = 4 r = 10 \(\dfrac{x}{360} = \dfrac{4}{2(10)\pi}\) \(x = \dfrac{360}{2(10) \pi}\)

OpenStudy (hrhsbull):

thanks